what is the imperical formula of an oxide of chromium is made up of 5.20 g chromium and 5.60 g oxygen?


1
Expert's answer
2021-10-05T01:45:49-0400

Moles of Chromium"=\frac{5.20}{52}=0.1"


Moles of oxygen "=\frac{5.60}{16}=0.35"


Mole ratio =(0.1÷0.1)=1 and

(0.35÷0.1)=3.5=4


"CrO_4"